OpenStudy (anonymous):

1. If a client puts the equivalent of $55 per month, or $660 per year, into an ordinary annuity, how much money would accumulate in 20 years at 3% compounded annually?

OpenStudy (anonymous):

396$. want me to teach you how to do it?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

Multiply 660 by 0.03, because if it says 3 percent, you divide it by 100, so it is 50%, divide 50 by 100. But for the problem, divide 660 by 0.03 because you need to find the interest per year, then multiply your answer by 20. Ok?

OpenStudy (sayan):

A=P(1+3/100)^t =660(1+0.03)^20 =1192.033

OpenStudy (anonymous):

thank you so much for your help

OpenStudy (sayan):

You Welcome...